OpenStudy (anonymous):

Reduce the following rational expression to lowest terms, if possible. (Problem in comments...) @ganeshie8

OpenStudy (anonymous):

\[\frac{ y+4 }{ y^2-16 }\]

OpenStudy (anonymous):

Hi ! @hartnn *waves* ^_^

hartnn (hartnn):

hello :)

OpenStudy (anonymous):

You two are thee best ! lol

ganeshie8 (ganeshie8):

I'll give one hint : \(a^2 - b^2 = (a+b)(a-b)\)

ganeshie8 (ganeshie8):

lol that motivates me for giving one more hint : \(16 = 4^2\)

OpenStudy (anonymous):

wouldnt the two y+4s cancel out? So ill just have y-4

ganeshie8 (ganeshie8):

You got it ! piece of cookie for u :)

hartnn (hartnn):

in the denominator, yes!

OpenStudy (anonymous):

lol, thank you...these are kinda easy...well so far lol dont go too far, you two! next question!

OpenStudy (anonymous):

Jeeze @hartnn and @ganeshie8 you guys are turning people into pros within 5 minutes. Teach me the ways!

OpenStudy (anonymous):

They are geniuses lol

ganeshie8 (ganeshie8):

too many kind words today lol xD ty :)

OpenStudy (anonymous):

wait, wait, wait! not done with the question !! Specify the restrictions on the variable. (Select all that apply.) y ≠ 16 y ≠ -4 y ≠ -16 y ≠ 4 y ≠ 0

ganeshie8 (ganeshie8):

you do knw that denominator can never become 0 right ?

OpenStudy (anonymous):

Yes, someone told me that

ganeshie8 (ganeshie8):

\(\dfrac{ y+4 }{ y^2-16 } = \dfrac{ y+4 }{ (y+4)(y-4) } = \dfrac{1 }{ y-4 } \)

ganeshie8 (ganeshie8):

to figure out the restrictions, look at ur `original` rational expression

ganeshie8 (ganeshie8):

\(\dfrac{ y+4 }{ y^2-16 } \)

ganeshie8 (ganeshie8):

you said, denominator cannot equal 0, so tell me what makes the denominator become 0 ?

ganeshie8 (ganeshie8):

those will be the restrictions

ganeshie8 (ganeshie8):

to find out the restrictions, you can simply set the denominator equal to 0 and solve : \(y^2 -16 = 0\) \(y = ?\)

OpenStudy (anonymous):

Tempted

OpenStudy (anonymous):

16? maybe 4 lol

ganeshie8 (ganeshie8):

\(y^2 -16 = 0 \\ y^2 = 16 \\ y = ?\)